OpenStudy (anonymous):

Ryder drew a scale drawing of the rectangular pool in his backyard using a scale of 1 in = 5 ft. He wants to add a border around the perimeter of the pool using 12-in square tiles. If the length of the drawing of the pool is 6 in, and the width of the drawing is 5 in, how many tiles does Ryder need for the border?

OpenStudy (ddcamp):

6 in on the drawing is 30 ft (6 in on page times 5 feet per inch) 5 inches on the drawing is 25 ft Perimeter = 2l + 2w = 2*30 + 2*25 = 60 + 50 =110 ft perimeter. 110 tiles around the edges, and then 4 extra (one for each corner) Grand total: 114 tiles

OpenStudy (anonymous):

thanks